OpenStudy (anonymous):

Put this in slope intercept form: 1. 2x-6y=6 2. 5x+3y=-15

OpenStudy (anonymous):

y=mx+b is the slope intercept. 1) -6y=-2x+6 and then you dived by 6 so you get y=1/3x-1

OpenStudy (anonymous):

thankx

OpenStudy (anonymous):

you gave me a medal thanks..lol

OpenStudy (anonymous):

np

OpenStudy (anonymous):

your welcome monica.